OpenStudy (anonymous):

what do y= for the equation 12y+6=6(2y+1)

OpenStudy (anonymous):

clear the brackets by distributing

OpenStudy (anonymous):

Ahhh, algebra. Whenever we have an equation solving for a variable, in this case 'y', we must isolate it. 12y+6=6(2y+1). I see y on both sides of the equation, but one of them is in a parentheses! Let's bust him out of there with distribution. 12y+6=12y+6. Strange. But let's continue. I want to subtract 6 from both sides to just have y. 12y=12y. Now let's divide by 12. y=y. Oh golly gee. The variable y equals itself! This means that y=R, where R is all real numbers.

OpenStudy (anonymous):

is this 1/2 of a 2 part equation?

OpenStudy (anonymous):

can the equation 12y+6=6(2y+ 1) be solved

OpenStudy (anonymous):

no

OpenStudy (anonymous):

y=y

OpenStudy (anonymous):

its the same as 2x = 2x x can be any number and the equation will always work

OpenStudy (anonymous):

yes I realize that y = all real numbers, but it can not be solved for a specific value.

OpenStudy (anonymous):

Isn't that what I said?

OpenStudy (anonymous):

yes you did a terrific job whatchalookinat :) but sometimes it seems not to sink in, but you deserve a medal